# Discovering universal temperature regulation dynamics in animals

### v1

@@ 1 Abstract
Hibernation is an adaptation to extreme environmental seasonality that has been studied for almost 200 years, but our mechanistic understanding of the underlying physiological system remains lacking due to the partially observed nature of the system. During hibernation, small mammals, such as the Arctic ground squirrel, exhibit dramatic oscillations in body temperature, typically one of the only physiological states measured, of up to 40 C. These spikes are known as interbout arousals and typically occur 10-20 times throughout hibernation. The physiological mechanism that drives interbout arousals is unknown, but two distinct mechanisms have been hypothesized. Using model selection for partially observed systems, we are able to differentiate between these two mechanistic hypotheses using only body temperature data recorded from a free-ranging Arctic ground squirrel. We then modify our discovered physiological model of Arctic ground squirrel to include environmental information and find that we can qualitatively match body temperature data recorded from a wide range of species, including a bird, a shrew, and a bear, which also dynamically modulate body temperature. Our results suggest that a universal, environmentally sensitive mechanism could regulate body temperature across a diverse range of species—a mechanistic restructuring of our current understanding of the physiological organization across species. While the findings presented here are applicable to thermophysiology, the general modeling procedure is applicable to time series data collected from partially observed biological, chemical, physical, mechanical, and cosmic systems for which the goal is to elucidate the underlying mechanism or control structure.
@@ 2 Introduction
Scientists have studied hibernation for nearly two hundred years [1] in the hopes of gaining mechanistic insight into this complex physiological system and to extend our own physiology. Hibernation research has a wide range of practical applications to human health, including cardiac arrest, organ preservation, and longevity. Inducing hibernation-like states into the range of human physiology is also under study, as this may help realize our species’ goals of long-distance human space travel [2, 3]. Interest in hibernation and dormancy states has transcended science and also fills our literature. Proust writes of “…the comfort of reclusion, the poetry of hibernation…” in A Swan’s Way. A more recent novel, My Year of Rest and Relaxation, follows a New York City woman who entered a months-long hibernation-like state. But for all this sustained scientific and broader interest, our mechanistic understanding of dynamic temperature regulation (known as the thermophysiological system) that facilitates hibernation remains limited, in part because few physiological states of the system can be observed experimentally. Typically, body temperature is the only physiological measurement that can be recorded from free-ranging organisms, though surely other states of the physiological system must contribute to thermoregulation. In applied mathematics, this is known as a partially observed system, and inferring models for this class of system is technically challenging and an active area of research [4, 5]. Here, we combine body temperature measurements with automated model discovery and classical dynamical systems theory to develop a mechanistic understanding of small mammalian thermophysiology and uncover a universal mechanism of thermophysiology across species.
@@
Relatively few mathematical models of hibernation and thermophysiology exist, and those that do exist tend to be either relatively complex and qualitative in nature or less mechanistic than is desired. Hampton and Andrews put forward a qualitative mechanistic model of small mammalian hibernation [10]. Their qualitative model puts forward a mechanism that triggers interbout arousals and can be viewed as a version of the Hourglass and Threshold Hypothesis. However, the model formulation is very high-dimensional and would be difficult to quantitatively validate using the available experimental data. The authors also developed a complex physiological model to infer metabolic rate from other physiological data [11]. Other less mechanistic frameworks have been used to examine features of hibernation. Oro and Freixas used statistical modeling to show that air temperature drives “flickering,” a behavior that occurs as the organism transitions between homeothermy (constant body temperature) and poikilothermy (spiking body temperature) [12]. A different study used Generalized Harmonic Analysis to investigate interbout arousals in small hibernators [13]. To differentiate between the Torpor Arousal Clock and Hourglass and Threshold Hypotheses, we use a recent model selection algorithm for partially observed systems to discover a simple, data-driven model of small mammalian hibernation. In our analysis, we recover the dynamics of a hidden physiological state, which has a constant period, lending support to the Hourglass and Threshold Hypothesis. We now use our discovered Arctic ground squirrel model to ask a broader inter-species question: Is our discovered thermophysiological mechanism universal across species with similar thermoregulatory behaviors?
@@
Switching between periods of homeothermy and poikilothermy is known as heterothermic endothermy and has been observed across many different species, including insects, small mammals, large mammals, and birds. It is thought to be an energy optimization strategy that is useful in a wide range of environments. Long-term body temperature data sets have been collected from these (and other) species, and the time series generally include some form of dramatic temporal oscillation [9, 14, 15, 16], but it is unclear how the patterns (and thus the underlying systems) relate to one another. Two different analyses attempted to answer this question. One analysis found that heterothermic endotherms can be broadly categorized into two (mostly) distinct groups: “daily heterotherms” and “hibernators,” though a few “intermediate” organisms fall somewhere between the two groups, including the elephant shrew [17, 18]. A different statistical analysis found that heterothermic endotherms may fall on a continuum [19, 20]. While valid, these statistical analyses are somewhat unsatisfying, as they do not tell us about the underlying thermoregulatory mechanisms and how that may or may not vary across species in a mechanistic sense. Here, we show that our selected physiological model of the Arctic ground squirrel, driven by environmental influences, such as photoperiod or temperature, can qualitatively capture body temperature patterns across species, including organisms in Geiser’s “daily heterotherm” category, “hibernator” category, and an “intermediate” organism [17]. This analysis uses a wide range of tools from applied mathematics to suggest a mechanistic hypothesis in thermal biology: a core thermophysiological mechanism is conserved across heterothermic endotherms, which can respond to and interact with environmental cues.
@@ 3.1 Model selection, robustness testing, and validation
To find a thermophysiological model for the Arctic ground squirrel, we employ a recent sparse model selection algorithm designed for partially observed systems [4]. In our system, only one variable, body temperature, is measured (Figure 2A). To determine the intrinsic dimensionality, we statistically analyze time-delay embeddings of body temperature data and find that the system is two-dimensional (Figure 2A) [21]. Thus, we seek to find a 2-D system of Ordinary Differential Equations (ODE) where the states are the measured temperature, $x$, and a hidden state, $y$, which represents unknown physiological dynamics. Sparse model selection evaluates a library of possible terms, ${\bf\Theta}(x,y)$, and composes ODE models by identifying a parsimonious set of terms that match observations of a subset of the system variables (Figure 2A-B). The search is performed by iteratively estimating the coefficients in the library using variational annealing and applying a threshold as a sparsity penalty. Iterative thresholding on the magnitude of the coefficients for each model term selects terms with large coefficients and sets small coefficients to zero in the next iteration. Of the models that emerge from varying sparsity and increasing model enforcement during optimization, those along the Pareto front balance simplicity and fit to the data, so we select the best fitting eight, nine, and ten active terms for further investigation (Figure 2B). [^1] For specific parameters the seven-term model exhibited a limit cycle, but cycling was not robust to small changes to any of the optimized parameters. Notice that these models form a nested family, and the algorithm selects variations on the same model. Our three models of interest can be written in the form
@@
To enable the models of interest to exhibit limit cycles with a finite period, we introduced a constant leak term, $\theta_{0}$, to the body temperature equation (see Equations (3)-(4)). We found that several low-order terms could be introduced to regulate the period of the limit cycle, but we chose to modify the data-driven models in the simplest way possible.  [^3] Notably, there exists a transformation from our models to Líenard equations when a linear term in $y$ is added to the body temperature equation [26]. With this simple addition to the models, the models of interest generate stable limit cycles with finite periods. To show that the limit cycle approaches an infinite period with zero leak, we introduced the leak term and simulated the Equations (3)-(4) (with $\theta_{C}=\theta_{D}=0$) as $\theta_{0}\rightarrow 0$. As expected, this resulted in limit cycle behavior with increasingly long periods (see Section 3.3 for technical explanation). Next, we perform additional robustness testing to determine which structural terms are necessary among the nested family of models.
@@ 3.2 Term robustness testing
Here, we test whether all remaining terms are necessary to produce the observed behavior based on overall term size. This is equivalent to keeping only terms that contribute significantly to the dynamics, consistent with dominant balance [27, 28, 29]. By forward simulating Equations (3)-(4) with $\theta_{C}=\theta_{D}=0$ and examining the term sizes of each component, we find that two additional terms, $\theta_{A}y^{3}$ and $\theta_{B}xy^{2}$, are negligible (see Figure 2C). By evaluating these terms across the body temperature time series, we see that these terms are relatively small compared to other terms (excluding $\theta_{0}$, which is needed to generate finite period limit cycle behavior). We now have a larger nested family of models of the form
@@
where the terms in black define the core model and the additional terms with coefficients, $\theta_{A},\theta_{B},\theta_{C},\theta_{D}$, are non-essential higher-complexity terms. Next, we examine how the model structure gives rise to limit cycle behavior.
@@
3.3 Qualitative behavior of the limit cycle and its parameter dependence
Here, we investigate how the oscillatory behavior emerges from a minimal model and depends on each of the dimensionless parameters. If we nondimensionalize Eqs. (3)-(4)) with $\theta_{A}$=$\theta_{B}$=$\theta_{C}$=$\theta_{D}$=0 and take the nondimensionalized form of $\theta_{4}$ ($\xi$) to be zero (see Section 6.3), we have
{dimensionless body $\{\text{dimensionless body temperature}\}~{}~{}\dot{\hat{x}}$ $=\hat{x}^{2}-\hat{x}^{3}-\gamma\hat{x}\hat{y}+\iota,\hat{x}(0)=\hat{x}_{0},$ (5)
{dimensionless unknown physiological $\{\text{dimensionless unknown physiological state}\}~{}~{}\dot{\hat{y}}$ $=\kappa\left(\hat{x}^{2}-\hat{y}^{2}\right),\hat{y}(0)=\hat{y}_{0}.$ (6)
This minimal system still exhibits limit cycle behavior, and each parameter is necessary to generate the behavior (see the limit cycle in phase space for $\gamma=1.2$, $\kappa=0.02$, and $\iota=$2 x 10-6 in Supplementary Figure 5). Note that when $\gamma=0$, $\dot{\hat{x}}$ no longer depends on $\hat{y}$, so $\hat{x}(t)$ will not have an oscillatory solution. If $\kappa$=0, $\hat{y}(t)$ is constant, and so the limit cycle is abolished by the same logic. The parameter $\iota$ is the key parameter that regulates the period length of the limit cycle. Notice that very close to the origin, the velocity in the phase plane in the $\hat{x}$-direction is almost completely dependent on $\iota$, as the quadratic and cubic terms become very small in this region. The velocity in the $\hat{y}$-direction will also be very small near the origin for the same reason. As $\iota\rightarrow 0$, the velocity in the $\hat{x}$ and $\hat{y}$ directions go to zero near the origin, elongating the period of the limit cycle. Therefore, there is an inverse relationship between $\iota$ and the period of the limit cycle, assuming $\iota<<1$.
@@ 3.4 Model validation
To quantitatively evaluate the fit of the model ensemble data to the data, we re-estimate the parameters of each model and compare against the training data (Figure 3A). We refer to the models by the number of active terms (not including the introduced leak term, $\theta_{0}$) and the subscript of the parameters of higher complexity terms the model contains (see Eqs. (3)-(4)). For example, Model 8AB refers to Eqs. (3)-(4) with $\theta_{A}\neq 0$, $\theta_{B}\neq 0$, and $\theta_{C}=\theta_{D}=0$. A complication in parameter estimation and validation for these models is that the interbout arousals of the Arctic ground squirrel are slightly aperiodic, perhaps due to small fluctuations in the decay rate of the unknown metabolite [8]. To generate a good fit, the forward simulation must match both the height of each body temperature spike and its timing, but a periodic spike will always be misaligned with the aperiodic data. To compensate, the peaks flatten to match most of the non-spiking data, which is near zero. To avoid forcing the data to match multiple periods, we fit all 16 models to the first two spikes of the training data (Figure 3A). We then validate the models on out-of-sample data using dynamic time warping.
@@
Which model is “best” depends on which quantities or qualities we measure. Many of our models are indistinguishable with the current dataset. If we examine the root mean-squared error between training data and forward simulation without penalizing for model complexity, the models are indistinguishable from each other based on the error (Figure 3A). If we penalize model complexity using the Akaike information criteria, the simplest Model 6 has the lowest score, but most of the seven-term models are statistically indistinguishable (Figure 3A, relative AIC $\leq 3$ dashed line). Dynamic time warping, which roughly compares the shape of the pulse corrected for time-shifting, supports Model 6, Model 7B, Model 7C, and 8BD (Figure 3A). If we consider more qualitative features of the fit, the simplest model, Model 6, has approximately the correct spike height, but spikes slightly too early and decays too fast (Figure 3B blue). More complex models, such as 9ABC (Figure 3B orange), introduce a refractory period after the interbout arousal, which the biological data does not exhibit. Determining which quantitative and subjective (shape of the fit) criteria to use for model selection is difficult without additional experimental information. For simplicity, we continue our analysis with Model 6, which is a reasonable fit to the data but not the only choice.
@@
Figure 3: Model validation shows that the simplest model, Model 6, is a reasonable fit to the data. A (top and middle): RMSE and AIC for the entire nested family of models fit to the first two interbout arousals of the training data. A (bottom): Dynamic time warping values of the model ensemble computed on the validation data. B (top): The fit of Model 6 (blue) to the training data (dotted purple). B (bottom): Close up of the fits during the first interbout arousal for Model 6 (blue) and Model 9ABC (orange), data show as dotted purple line. C (top): Dynamics of Model 6 ($x(t)$ in blue, $y(t)$ in pink). The hidden state has a fixed period, lending support to the Hourglass and Threshold Hypothesis [8]. C (bottom): Introducing a scaling symmetry into the hidden state changes the biological interpretation. In this case, the physiological threshold acts on the difference between multiple metabolites rather than a single metabolite.

### v2

@@ 1 Abstract
Hibernation is an adaptation to extreme environmental seasonality that has been studied for almost 200 years, but our understanding of the underlying physiological system remains lacking due to the partially observed nature of the system. During hibernation, small mammals, such as the Arctic ground squirrel, exhibit dramatic oscillations in body temperature, typically one of the only physiological states measured, of up to 40 C. These spikes are known as interbout arousals and typically occur 10-20 times throughout hibernation. The physiological process that drives interbout arousals is unknown, but two distinct macro-scale mechanisms have been hypothesized. Using model selection for partially observed systems and classical dynamical systems theory, we are able to differentiate between these two hypotheses using only body temperature data recorded from a free-ranging Arctic ground squirrel, and show that our model can capture the broad features of the observed seasonal physiological transitions. We then modify our discovered physiological model of Arctic ground squirrel to include environmental information and find that we can qualitatively match body temperature data recorded from a wide range of species, including a bird, a shrew, and a bear, which also dynamically modulate body temperature. Our results suggest that a universal, environmentally sensitive core regulator could control body temperature across a diverse range of species—a restructuring of our current understanding of the physiological organization across species. While the findings presented here are applicable to thermophysiology, the general modeling procedure is applicable to time series data collected from partially observed biological, chemical, physical, mechanical, and cosmic systems for which the goal is to elucidate the underlying mechanism or control structure.
@@ 2 Introduction
Scientists have studied hibernation for nearly two hundred years [1] in the hopes of gaining insight into this complex physiological system and to extend our own physiology. Hibernation research has a wide range of practical applications to human health, including cardiac arrest, organ preservation, and longevity. Inducing hibernation-like states into the range of human physiology is also under study, as this may help realize our species’ goals of long-distance human space travel [2, 3]. Interest in hibernation and dormancy states has transcended science and also fills our literature. Proust writes of “…the comfort of reclusion, the poetry of hibernation…” in A Swan’s Way. A more recent novel, My Year of Rest and Relaxation, follows a New York City woman who entered a months-long hibernation-like state. But for all this sustained scientific and broader interest, our understanding of dynamic temperature regulation (known as the thermophysiological system) that facilitates hibernation remains limited, in part because few physiological states of the system can be observed experimentally. Typically, body temperature is the only physiological measurement that can be recorded from free-ranging organisms, though surely other states of the physiological system must contribute to thermoregulation. In applied mathematics, this is known as a partially observed system, and inferring models for this class of system is technically challenging and an active area of research [4, 5]. Here, we combine body temperature measurements with automated model discovery and classical dynamical systems theory to develop a mechanistic understanding of small mammalian thermophysiology and uncover a universal mechanism of thermophysiology across species.
@@
Relatively few mathematical models of hibernation and thermophysiology exist, and those that do exist tend to be either relatively complex and thus difficult to validate against the existing data or statistical in nature. Hampton and Andrews put forward a mechanistic model of small mammalian hibernation [10]. Their model puts forward a mechanism that triggers interbout arousals and can be viewed as a version of the Hourglass and Threshold Hypothesis. However, the model formulation is very high-dimensional and cannot be validated using the available experimental data. The authors also developed a complex physiological model to infer metabolic rate from other physiological data [11]. Others have used statistical frameworks to examine features of hibernation. Oro and Freixas used statistical modeling to show that air temperature drives “flickering,” a behavior that occurs as the organism transitions between homeothermy (constant body temperature) and poikilothermy (spiking body temperature) [12]. A different study used Generalized Harmonic Analysis to investigate interbout arousals in small hibernators [13]. We seek to find the right level of mathematical description for hibernation physiology that balances interpretability and preserves the ability to validate the model against the existing data. This is challenging due to natural and experimental limitations intrinsic to the system.
Hibernation is a multi-scale physiological adaptation that occurs on the year scale. For this reason, a molecular description of the biological processes that underpin hibernation will be exceedingly difficult to find. It is unlikely we will ever have complete and comprehensive data sets of all relevant neurophysiological subsystems involved in hibernation due to the costs of studying hibernators from both a financial and animal welfare perspective. Most of the data collected from hibernation experiments occurs at the physiological scale rather than the molecular scale. We grappled with the right size and type of model to use in describing this system, and ultimately found a model that is roughly as mechanistic the Fitzhugh-Nagamo model [14, 15], which describes neuronal spiking dynamics; the NPZ model [16], which describes the dynamics of nutrients, phytoplankton and zooplankton; or the Brusselator model [17, 18, 19] which describes the Belousov-Zhabotinsky chemical reaction [20]. The goal of these models, and perhaps all models, is to describe the essence of the underlying process in an interpretable manner without being completely overrun by the complexity of the system or the curse of dimensionality. Balancing model simplicity, which allows for interpretability, with complexity, which allows for greater realism, is a fine line. The NPZ model is a good example of a model that exhibits this balance. It captures a complex, multi-scale ecological system using just three coupled equations, and can be used to probe how changes or perturbations to one state influence the others, which would be challenging to understand both without any model and with too complex of a model. Here, we attempt to find a similar modeling balance. We discover a macro-scale model with interpretable parameters that have physiological meaning, that is complex enough to capture the main features of the data, that is simple enough to be fit to the existing physiological data, and can be used to guide the search for microscale mechanisms that underpin hibernation. Using our discovered model, we provide new insights into hibernation physiology, seasonal physiological transitions, and the possible conserved physiological organization across species.
To find this model, we use a recent sparse model selection algorithm for partially observed systems and discover a simple, data-driven model of small mammalian hibernation. We then use this model to differentiate between the Torpor Arousal Clock and Hourglass and Threshold Hypotheses. In our analysis, we recover the dynamics of a hidden physiological state, which has a constant period, lending support to the Hourglass and Threshold Hypothesis. We also use our discovered model to explore seasonal thermoregulatory transitions through a bifurcation analysis, which results in a multi-scale understanding of hibernation that links the neurobiological and physiological scales. As a final test of our discovered Arctic ground squirrel model, we used it to explore a broader, interspecific question: Could a core thermophysiological controller be responsible for regulating similar thermoregulatory behaviors across species?
@@
During the annual cycle, the Arctic ground squirrel switches between two distinct thermoregulatory behaviors: a stable, regulated body temperature and a dynamically changing body temperature marked by temporal oscillation. Many other organisms exhibit similar transitions between the two thermoregulatory modes, but the timescale of the behaviors and timing of transitions between them is different. Long-term body temperature data sets have been collected from these (and other) species [9, 21, 22, 23], but it is unclear how the patterns, and thus the underlying systems, relate to one another. A satisfactory model would be capable of capturing a broad range of observed body temperature patterns across species and would further our understanding of physiological organization across species. Previous analyses attempted to categorize body temperature recordings across species using a statistical lens rather than applying a dynamical systems interpretation. One analysis found that heterothermic endotherms can be broadly categorized into two (mostly) distinct groups: “daily heterotherms” and “hibernators,” though a few “intermediate” organisms fall somewhere between the two groups, including the elephant shrew [24, 25]. A different statistical analysis found that heterothermic endotherms may fall on a continuum [26, 27]. Here, we show that our selected physiological model of the Arctic ground squirrel, driven by environmental influences, such as photoperiod or temperature, can qualitatively capture body temperature patterns across species, including organisms in Geiser’s “daily heterotherm” category, “hibernator” category, and an “intermediate” organism [24]. This analysis uses a wide range of tools from applied mathematics to suggest a mechanistic hypothesis in thermal biology: a core thermophysiological mechanism is conserved across heterothermic endotherms, which can respond to and interact with environmental cues.
@@ 3.1 Model selection, robustness testing, and validation
To find a thermophysiological model for the Arctic ground squirrel, we employ a recent sparse model selection algorithm designed for partially observed systems [4]. In our system, only one variable, body temperature, is measured (Figure 2A). To determine the intrinsic dimensionality, we statistically analyze time-delay embeddings of body temperature data and find that the system is two-dimensional (Figure 2A and Section 6.2) [28]. Thus, we seek to find a 2-D system of Ordinary Differential Equations (ODE) comprised of monomial terms where the states are the measured temperature, $x$, and a hidden state, $y$, which represents the dynamics of an unknown physiological process. This allows us to find an effective macro-scale physiological model describing the dynamics of coupling between temperature and some unknown internal control process. Sparse model selection evaluates a library of possible terms, ${\bf\Theta}(x,y)$, and composes ODE models by identifying a parsimonious set of terms that match observations of a subset of the system variables (Figure 2A-B). The search is performed by iteratively estimating the coefficients in the library using variational annealing and applying a threshold as a sparsity penalty. Iterative thresholding on the magnitude of the coefficients for each model term selects terms with large coefficients and sets small coefficients to zero in the next iteration. Of the models that emerge from varying sparsity and increasing model enforcement during optimization, those along the Pareto front balance simplicity and fit to the data, so we select the best fitting eight, nine, and ten active terms for further investigation (Figure 2B). [^1] For specific parameters the seven-term model exhibited a limit cycle, but cycling was not robust to small changes to any of the optimized parameters. These models form a nested family, and the algorithm selects variations on the same model. Our three models of interest can be written in the form
@@
To enable the models of interest to exhibit limit cycles with a finite period, we introduced a constant leak term, $\theta_{0}$, to the body temperature equation (see Equations (3)-(4)). We found that several low-order terms could be introduced to regulate the period of the limit cycle, but we chose to modify the data-driven models in the simplest way possible.  [^3] Notably, there exists a transformation from our models to Líenard equations when a linear term in $y$ is added to the body temperature equation [31]. With this simple addition, the models of interest generate stable limit cycles with finite periods. To show that the limit cycle approaches an infinite period with zero leak, we introduced the leak term and simulated the Equations (3)-(4) (with $\theta_{C}=\theta_{D}=0$) as $\theta_{0}\rightarrow 0$. As expected, this resulted in limit cycle behavior with increasingly long periods, as the phase plane velocity near the origin goes to zero as $\theta_{0}\rightarrow 0$. Next, we perform additional robustness testing to determine which structural terms are necessary among the nested family of models.
@@ 3.2 Term robustness testing
Here, we test whether the remaining terms are necessary to produce the observed behavior based on overall term size. This is equivalent to keeping only terms that contribute significantly to the dynamics, consistent with dominant balance [32, 33, 34]. By forward simulating Equations (3)-(4) with $\theta_{C}=\theta_{D}=0$ and examining the term sizes of each component, we find that two additional terms, $\theta_{A}y^{3}$ and $\theta_{B}xy^{2}$, are negligible across the time-series of a spike (see Figure 2C). The terms $\theta_{0}$, $\theta_{5}x^{2}$, and $\theta_{6}y^{2}$ are excluded from this analysis as they are required to generate finite period limit cycle behavior. We now have a larger nested family of models of the form
@@
where the terms with coefficients $\theta_{1},\theta_{2},\theta_{3},\theta_{4},\theta_{5},\theta_{6}$ (in black), define the core model and the additional terms with coefficients, $\theta_{A},\theta_{B},\theta_{C},\theta_{D}$ (blue and pink, Figure 2 D), are non-essential higher-complexity terms. Next, we validate our models using the $l^{2}$ norm, information criteria, and dynamic time warping.
@@ 3.3 Model validation
To quantitatively evaluate the fit of the model ensemble data to the data, we re-estimate the parameters of each model and compare against the training data (Figure 3A). We refer to the models by the number of active terms (not including the introduced leak term, $\theta_{0}$) and the subscript of the parameters of higher complexity terms that the model contains (see Eqs. (3)-(4)). For example, Model 8AB refers to Eqs. (3)-(4) with $\theta_{A}\neq 0$, $\theta_{B}\neq 0$, and $\theta_{C}=\theta_{D}=0$. A complication in parameter estimation and validation for these models is that the interbout arousals of the Arctic ground squirrel are slightly aperiodic, perhaps due to small fluctuations in the decay rate of the unknown metabolite [8]. To generate a good fit, the forward simulation must match both the height of each body temperature spike and its timing, but a periodic spike will always be misaligned with the aperiodic data. To compensate, the peaks flatten to match most of the non-spiking data, which is near zero. To avoid forcing the data to match multiple periods, we fit all 16 models to the first two spikes of the training data (Figure 3A). We then evaluate the fit of the models using the $l^{2}$ norm and information criteria on the first two spikes of the training data, and validate the models on out-of-sample data using dynamic time warping.
@@
Which model is “best” depends on which quantities or qualities we measure. Many of our models are indistinguishable with the current dataset. If we examine the $l^{2}$ norm between training data and forward simulation without penalizing for model complexity, the models are indistinguishable from each other based on the error (Figure 3A). If we penalize model complexity using the Akaike information criteria, the simplest Model 6 has the lowest score based on the training data, but most of the seven-term models are statistically indistinguishable (Figure 3A, relative AIC $\leq 3$ dashed line). Applying dynamic time warping to the out-of-sample data, which roughly compares the shape of the pulse corrected for time-shifting, supports Model 6, Model 7B, Model 7C, and 8BD (Figure 3A). If we consider more qualitative features of the fit, the simplest model, Model 6, has approximately the correct spike height, but spikes slightly too early and decays too fast (Figure 3B blue). More complex models, such as 9ABC (Figure 3B orange), introduce a refractory period after the interbout arousal, which the biological data does not exhibit. Determining which quantitative and subjective (shape of the fit) criteria to use for model selection is difficult without additional experimental information. For simplicity, we continue our analysis with Model 6, which is a reasonable fit to the data but not the only choice.
@@
Figure 3: Model validation shows that the simplest model, Model 6, is a reasonable fit to the data. A (top and middle): $l^{2}$ norm and AIC for the entire nested family of models fit to the first two interbout arousals of the training data. A (bottom): Dynamic time warping values of the model ensemble computed on the validation data. B (top): The fit of Model 6 (blue) to the training data (dotted purple). B (bottom): Close up of the fits during the first interbout arousal for Model 6 (blue) and Model 9ABC (orange), data shown as a dotted purple line. C (top): Dynamics of Model 6 ($x(t)$ in blue, $y(t)$ in pink). The hidden state has a fixed period, lending support to the Hourglass and Threshold Hypothesis [8]. C (bottom): Introducing a scaling symmetry into the hidden state changes the biological interpretation. In this case, the physiological threshold acts on the difference between multiple metabolites rather than a single metabolite.